OpenStudy (anonymous):

2-xz=0 -4yz-3=0 -x^2-2y^2+4=0 solve for x and y

OpenStudy (diyadiya):

xz=2 so x =2/z 4yz = -3 so y= -3/4z substitute it in equation x^2 + 2y^2 = 4 and find z then similarly find x and y